ter Problems 1. Use finite differences to approximate solutions to the linear BVPs for n=9, 19, and 39. (a) | y = y + ſet {MATH LAB CODE PLS
Answer #1

Dear Student,

Following is the answer:

(a) Consider the following linear Boundary Value Problem, y(0) = 0 »(1) =- The above differential equation at 1, in the discr

Here, ya and by are the boundary conditions.

Differentiate y, -(2+h?) w; -25% e + W, = 0) with respect to w; . Thus, -(2+h+)=0 y(0) = 0, y(1) --- Use the following MATLA

%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%% function y=f(Winter, bv,n) y=zeros(n,1); h=(inter(2)-inter(1)/(n+1); y(1)=bv(1)-(2+h^2)*

HOME PLOTS APPS Ayon 2 County New New Open compare port Des We w e - - - - Users De MATLAB >> veloved to 1). 10 exp (1)/). 20

The resulting solution curve yt is :

CHU Times New Roman - 12 - File Edit View the Tool Detop Window Help AaB obce. A nd S . S Approximate solutions of BVP by the

The exact solution for the linear BVP Y = y + e,y(0) = 0, y(1)=-e is given by, v(t) = zce Use the following MATLAB program

for i=0:10 t(i+1)=t(i+1)+i/10; z(i+1)=z(i+1)+t(i+1)*exp(t(i+1))/3; end plot([a a+(1:n)*h b].[ya W yb]);hold on plot(t,z) xlaa(i,i+1)=1; a (i+1,1)=1; end Use the MATLAB command w=nlbvpfdia( [0 1],[0 exp(1)/3],9) to plot the approximate solutions toMARA 2013 HOME PLOTS APPS w e Es www Open de View that too Desktop Window DOBRODD Plot of the approximate and the exact solut

Use the following MATLAB program to calculate errors and make a semilog plot for the errors. This program is saved as nlbvpf

z(i)=z(i)+t(i)*exp(t1))/3; err(i)=err(i)+abs(w(1)-z(i)); end disp([ Errors=]) disp(err) semilogy(t.err) xlabel(t),ylabel(

end for i=1:n-1 a(i,i+1)=1; a (i+1,1)=1; end Use the MATLAB command w=nlbvpfdia( [01],[O exp(1)/3],9) to calculate errors a

MAR 2018 HOME PLOTS APPS s e Save New Come t de de view met Tool Desktop Window Help waa94. GOED A semilog plot for the error

(b) Consider the following linear Boundary Value Problem, y = (2+41?)y y(0)=1 y(1)=e The above differential equation at 1, i

Here, ya and by are the boundary conditions.

Use the following MATLAB program to approximate solutions for the linear BVP y = (2+41° ) y, y(0)=1, y(1)=e. This program is

y(1)=bv(1)-(2+2+h^2+4*h^4)*W(1)+W(2); y(n)=W(n-1)-(2+2*h^2+4**40*W(n)+bv(2); for i=2:n-1 y(i)=W(i-1)-(2+2*h^2+4*h^4)*W(1)+W(i

Consider n=9. Use the MATLAB command w=nlbvpfdib( [01],[1 exp(1)],9) to approximate solutions for the linear BVP y = (2+41

HOME APPS L PLOTS E compare Wate Aye Cute h oto 2 Community New New Open port Gew e - Ch . The Users De MATLAB 19EPAT I Error

The resulting solution curve y(t) is given below:

MARA 2013 HOME PLOTS APPS F porte e e t View It To Desktop Window Help New New Open compare Approximate solutions of BVP by t

The exact solution for the linear BVP y = (2+41?)y, y(0) = 1, y(1)= e is given by, y(t)=en Use the following MATLAB program

for i=0:10 t(i+1)=t(i+1)+i/10; z(i+1)=z(i+1)+exp(t(i+1)^2); end disp([The exact solutions are]) disp(z) plot([a a+(1:n)*h bfor i=1:n ali,i)=-2-2*h^2-4*h^4; end for i=1:n-1 a(i,i+1)=1; a (i+1,1)=1; end Use the MATLAB command w=nlbvpfdib( [01],[1 exONU File Edit View AaB BCAA and Tool Detop Window Help OR GOOD Plot of the approximate and the exact solutions AaB 1 e Same S

t(i)=t(i)+i/10; z(i)=z(i)+t(i)*exp(ti))/3; err(i)=err(i)+abs(w(1)-z(i)); end disp(Errors=]) disp(err) semilogy(t,err) xlabe

a(i,i)=-2-2*h^2-4*h^4; end for i=1:n-1 a(i,i+1)=1; a (i+1,1)=1; endUse the MATLAB command w=nlbvpfd1b2( [01],[1 exp(1)],9) to calculate errors and make a semilog plot for the errors. See theHU ... 1 AN File Edit View the Tool Desktop Window Help Times New Roman AaB chce. A nd S e Suite Change Select A semilog plot

Consider n=19 Use the MATLAB command w=nlbvpfdib( [01],[1 exp(1)],19) to approximate solutions for the linear BVP b) = 1, y(

ONU File Edit View the Tool Desktop Window Help AdD ABG ABOCA Approximate solutions of BVP by the Finite Differrace Method Ap

Consider n = 39 Use the MATLAB command w=nlbvpfdi( [01],[1 exp(1)],39) to approximate solutions for the linear BVP y = (2+

MARA 2013 HOME PLOTS APPS F www com Nome mot Save e e t View It To Desktop Window Help waaASSO926.& DEOD Approximate solution